if 65 is added to a number, the result is 39 less than three times the number. What is the number?

Answer by checkley79(3341)

X+65=3X-39
X-3X=-39-65
-2X=-104
X=-104/-2
X=52 ANS.
PROOF:
52+65=3*52-39
117=156-39
117=117